a piece of metal(mass=18.300g) is placed in 11.00ml of chloroform(d=1.498g/ml) in a 25ml graduated cylinder. The chloroform level increases to 15.46. the best value for density of metal from the data
1
Expert's answer
2020-09-21T06:46:23-0400

The density of the piece of metal can be calculated as following:

d = m/V

where m - mass, V - volume.

The volume of the piece of metal equals the difference between the initial volume of chloroform and the final volume of chloroform in the graduated cylinder. As a result:

V(metal) = V(chloroform final) - V(chloroform initial) = 15.46 ml - 11.00 ml = 4.46 ml

From here:

d(metal) = m(metal) / V(metal) = 18.300 g / 4.46 ml = 4.10 g/ml


Answer: 4.10 g/ml